180 x + 52 = 26 + 90x
Answer by waynest(281)

180x + 52 = 26 + 90x
180x - 90x + 52 - 52 = 26 - 52 + 90x - 90x
90x = -26
90x/90 = -26/90
x = -13/45
check:
180(-13/45) + 52 = 26 + 90(-13/45)
-52 + 52 = 26 - 26
0 = 0